# The Spectral Edge Thesis: A Mathematical Framework for Intra-Signal Phase Transitions in Neural Network Training

### 30 Mar 2026 v1

@@ 41 Circuit Survival and the Final Model
A pretrained GPT-2 has many identifiable circuits (induction heads, IOI circuits, etc.), yet we observe $k^{*}=2$33 and only $\sim 2$ spectral edge events in the observation window. The resolution is the holographic encoding (Section 39): each event is rank-1 in function space but maps through $\bm{J}^{\dagger}$ to a dense reorganisation of the entire parameter space. Mechanistic interpretability decomposes this one dense change into multiple “circuits” because it analyses the computational graph, not function space. The complexity of the final model comes from the dimensionality of parameter space ($P\sim 10^{8}$), not from the number of spectral edge events. Additionally, many events occurred before the observation window—the Gram matrix is a local diagnostic, not a historical record.
@@ 42 The Geometric Flow Connection
Preliminary evidence from the Muon experiment (Open Problem 8 below) is suggestive: AdamW and Muon produce different spectral edge dynamics ($k^{*}=2$ vs. $k^{*}=1$, $R\approx 10$$30$ vs. $R\approx 3$55) yet reach models with comparable capabilities. The flow differs; the invariant appears preserved.
@@ 43 Logical Structure and the Commutativity Assumption
For SGD ($\mathcal{P}=I$), commutativity is exact and all results hold. For Adam with $\beta_{2}\to 1$, it is approximately satisfied. For Muon, it is the most questionable—see Open Problem 8 below.
@@ 44 Open Problems
    Scaling of $k^{*}$ with Model Size and Optimizer. Our empirical data shows $k^{*}\in\{2,3\}$ for models of size 51M–124M under AdamW. The Muon experiment (Section 44) provides direct evidence that $k^{*}$ is at least partly optimizer-dependent: Muon’s spectral norm equalisation drives $k^{*}=1$, whereas AdamW produces $k^{*}=2$ on the same TinyStories 51M model, while both reach comparable validation loss. Whether $k^{*}$ remains small at billion-parameter scale, and how it varies across optimizers and architectures, are open empirical questions. The Krylov bound (Proposition 5.5) suggests $k^{*}$ depends on the number of Hessian outliers rather than $p$ directly, but this has not been tested beyond 124M parameters.
@@
    Preconditioned optimisers: the Muon test. The Hessian–trajectory gap correspondence (Prop. 5.1) requires commutativity $\mathcal{P},H]\approx 0$. For SGD this is exact; for Adam it is approximate. The Muon optimiser (momentum + orthogonalisation via Newton–Schulz [31]) has a nonlinear, state-dependent preconditioner $P(m)=(mm^{\top})^{-1/2}$ that equalises spectral norms, partially counteracting the Hessian’s role in creating the gap.

### 03 Apr 2026 v2

@@ 41 Circuit Survival and the Final Model
A pretrained GPT-2 has many identifiable circuits (induction heads [28], IOI circuits [29], etc.), yet we observe $k^{*}=2$33 and only $\sim 2$ spectral edge events in the observation window. The resolution is the holographic encoding (Section 39): each event is rank-1 in function space but maps through $\bm{J}^{\dagger}$ to a dense reorganisation of the entire parameter space. Mechanistic interpretability decomposes this one dense change into multiple “circuits” because it analyses the computational graph, not function space. The complexity of the final model comes from the dimensionality of parameter space ($P\sim 10^{8}$), not from the number of spectral edge events. Additionally, many events occurred before the observation window—the Gram matrix is a local diagnostic, not a historical record.
@@ 42 The Geometric Flow Connection
Preliminary evidence from the Muon experiment (Open Problem 9 below) is suggestive: AdamW and Muon produce different spectral edge dynamics ($k^{*}=2$ vs. $k^{*}=1$, $R\approx 10$$30$ vs. $R\approx 3$55) yet reach models with comparable capabilities. The flow differs; the invariant appears preserved.
@@ 43 Logical Structure and the Commutativity Assumption
For SGD ($\mathcal{P}=I$), commutativity is exact and all results hold. For Adam with $\beta_{2}\to 1$, it is approximately satisfied. For Muon, it is the most questionable—see Open Problem 9 below.
@@ 44 Open Problems
    Scaling of $k^{*}$ with Model Size and Optimizer. Our empirical data shows $k^{*}\in\{2,3\}$ for models of size 51M–124M under AdamW. The Muon experiment (Open Problem 9) provides direct evidence that $k^{*}$ is at least partly optimizer-dependent: Muon’s spectral norm equalisation drives $k^{*}=1$, whereas AdamW produces $k^{*}=2$ on the same TinyStories 51M model, while both reach comparable validation loss. Whether $k^{*}$ remains small at billion-parameter scale, and how it varies across optimizers and architectures, are open empirical questions. The Krylov bound (Proposition 5.5) suggests $k^{*}$ depends on the number of Hessian outliers rather than $p$ directly, but this has not been tested beyond 124M parameters.
@@
    Preconditioned optimisers: the Muon test. The Hessian–trajectory gap correspondence (Proposition 5.1) requires commutativity $\mathcal{P},H]\approx 0$. For SGD this is exact; for Adam it is approximate. The Muon optimiser (momentum + orthogonalisation via Newton–Schulz [31]) has a nonlinear, state-dependent preconditioner $P(m)=(mm^{\top})^{-1/2}$ that equalises spectral norms, partially counteracting the Hessian’s role in creating the gap.